find the minimum value of y=x2-22x+132
Answer by stanbon(75887)

find the minimum value of y=x^2-22x+132
----
Minimum occurs when x = -b/(2a) = 22/(2) = 11
f(11) = 11^2-22*11+132 = 121-242+132 = 11
============
Minimum: (11,11)
================
Cheers,
Stan H.
===============